OpenStudy (anonymous):

Four expressions are shown below: 2(4x + 2) 2(3x + 3) 6x + 6 8x + 4 Which two expressions are equivalent to 2(3x + 2 + x)? A and D B and D A and C C and B @Holly00d1248 @larryboxaplenty @dtan5457

OpenStudy (dtan5457):

Multiply the answer out first 2(3x + 2 + x)=6x+4+2x=8x+4 Which of your answers = 8x+4?

OpenStudy (anonymous):

B and D ? @dtan5457

OpenStudy (holly00d1248):

nope if you combine lie terms 3x and 1x =4x now wat you think

OpenStudy (holly00d1248):

*like

OpenStudy (dtan5457):

D is one of them, but B is not.

OpenStudy (anonymous):

d an A? @dtan5457

OpenStudy (holly00d1248):

yeps

OpenStudy (dtan5457):

Yes.

OpenStudy (michele_laino):

what is: 3x + 2 + x=...?

OpenStudy (holly00d1248):

4x+2 why